# Recoloring via modular decomposition

### v1

@@ Abstract
The reconfiguration graph of the $k$-colorings of a graph $G$, denoted $R_{k}(G)$, is the graph whose vertices are the $k$-colorings of $G$ and two colorings are adjacent in $R_{k}(G)$ if they differ in color on exactly one vertex. A graph $G$ is said to be recolorable if $R_{\ell}(G)$ is connected for all $\ell\geq\chi(G)$+1. We use the modular decomposition of several graph classes to prove that the graphs in the class are recolorable. In particular, we prove that every ($P_{5}$, diamond)-free graph, every ($P_{5}$, house, bull)-free graph, and every ($P_{5}$, $C_{5}$, co-fork)-free graph is recolorable.
@@ 1 Introduction
Let $G$ be a finite simple graph with vertex-set $V(G)$ and edge-set $E(G)$. We assume that $G$ is connected unless stated otherwise. We use $n$ to represent the number of vertices in a graph. Two vertices $u$ and $v$ are adjacent in $G$ if $uv\in E(G)$. For a positive integer $k$, a $k$-coloring of $G$ is a mapping from $V(G)$ to a set of colors $\{1,2,\dots,k\}$ such that no two adjacent vertices receive the same color. We say that $G$ is $k$-colorable if it admits a $k$-coloring, and the chromatic number of $G$, denoted $\chi(G)$, is the minimum number of colors required to color $G$. We say $\chi$-coloring of $G$ instead of $\chi(G)$-coloring of $G$, when appropriate. The reconfiguration graph of the $k$-colorings, denoted $R_{k}(G)$, is the graph whose vertices are the $k$-colorings of $G$ and two colorings are adjacent in $R_{k}(G)$ if they differ in color on exactly one vertex. A graph $G$ is said to be $k$-mixing if $R_{k}(G)$ is connected. Given a reconfiguration graph, we can ask the following questions: Is the reconfiguration graph connected? If so, what is the diameter of the reconfiguration graph? In this paper we focus on the connectivity of the reconfiguration graph of the $k$-colorings. We say a graph $G$ is recolorable if $R_{\ell}(G)$ is connected for all $\ell\geq\chi(G)$+1.
@@
Deciding whether there exists a path between two colorings in $R_{k}(G)$ is PSPACE-complete for $k>3$ [6], and can be decided in polynomial time for $k\leq 3$ [8]. In 2018, Wrochna proved that the problem remains PSPACE-complete for graph classes with bounded bandwidth, and hence for graph classes with bounded treewidth [18]. This led researchers to study the problem for various restricted graph classes. In [5], a graph class known as $k$-color-dense graphs was introduced and it was proved that every $k$-color-dense graph is recolorable. The class of $k$-color-dense graphs includes the class of $k$-colorable chordal graphs. For a graph $H$, a graph $G$ is $H$-$free$ if no induced subgraph of $G$ is isomorphic to $H$. In [2], it was proved that every $H$-free graph $G$ is recolorable if and only if $H$ is an induced subgraph of $P_{4}$ or $P_{3}$+$P_{1}$. For a collection of graphs $\mathcal{H}$, $G$ is $\mathcal{H}$-free if $G$ is $H$-free for every $H\in\mathcal{H}$.
@@
Bonamy and Bousquet [4] used the tree decomposition of graphs to show that every graph $G$ is ($tw(G)$+2)-mixing, where $tw(G)$ denotes the treewidth of the graph $G$. In this paper we use the modular decomposition of a graph $G$ to prove that $G$ is recolorable.
@@
A set $S\subseteq V(G)$ is a module if $1\leq|S|\leq|V(G)|$ and every vertex in $V(G)\setminus S$ is either adjacent to every vertex of $S$ or no vertex of $S$. A module $S$ is said to be non-trivial if $S$ contains at least two vertices. Note that for each $v\in V(G)$, $\{v\}$ is a trivial module in $G$. A graph $G$ is prime if it does not contain any non-trivial module.
@@
A class of graphs $\mathcal{G}$ is said to be hereditary if for every $G$ in $\mathcal{G}$, every induced subgraph of $G$ is also in $\mathcal{G}$. Note that for a collection of graphs $\mathcal{H}$, the class of $\mathcal{H}$-free graphs is hereditary. In Section 3, we prove the following.
@@
Let $\mathcal{G}$ be a hereditary class of graphs and let $\mathcal{H}$ be the set of prime graphs in $\mathcal{G}$. If for all $H\in\mathcal{H}$, every blowup of $H$ is recolorable, then for all $G\in\mathcal{G}$, $G$ is recolorable.
We use the above result in Section 4 along with the modular decomposition of several subclasses of $P_{5}$-free graphs to prove the following.
@@ 3 Modular Decomposition
For the remainder of this section, we use the following decomposition. Let $G$ be a graph which is neither a join nor a disjoint union of two graphs. Let $S_{1}$,…, $S_{m}$ be a partition of $V(G)$ into maximal modules. Note that $S_{p}\cap S_{q}$ = $\emptyset$ for all $p\neq q$, $p$ and $q\in[m]$. Let $\chi(G)=k$ and let $\chi(G[S_{p}])=k_{p}$, for all $p\in[m]$. For a graph $G$, the clique skeleton of $G$ is the graph obtained from $G$ by recursively substituting a clique $Q_{p}$ of size $k_{p}$ for each $S_{p}$ in $G$, where $p\in[m]$. We use $G(Q_{1},\dots,Q_{m})$ to denote the clique skeleton of $G$. Similar operations have been defined in the literature, for example see [15]. Let $H$ = $G(Q_{1},\dots,Q_{m})$ be the clique skeleton of $G$. Then for any $k$-coloring $\alpha$ of $G$, there exists a $k$-coloring $\beta$ of $H$ such that $\beta(Q_{p})\subseteq\alpha(S_{p})$, for all $p\in[m]$. Note that such a coloring $\beta$ of $H$ can be obtained by coloring each $Q_{p}$ with colors from the set $\alpha(S_{p})$, for all $p\in[m]$. Similarly, given any $k$-coloring $\beta$ of $H$, there exists a $k$-coloring $\alpha$ of $G$ such that for all $p\in[m]$, $\alpha(S_{p})$ = $\beta(Q_{p})$. Thus $\chi(G)$ = $\chi(H)$. Note that the clique skeleton of a graph $G$ is a blowup of the skeleton of $G$.
@@
Let $\mathcal{G}$ be a hereditary class of graphs and let $\mathcal{H}$ be the set of prime graphs in $\mathcal{G}$. Assume that for all $H\in\mathcal{H}$, every blowup of $H$ is recolorable. If possible, let $G\in\mathcal{G}$ be a vertex-minimal graph that is not recolorable. Then $G$ is neither a join nor a disjoint union of two graphs. Let $S_{1}$,…, $S_{m}$ be a partition of $V(G)$ into maximal modules. Let $G^{*}$ be the skeleton of $G$. Then $G^{*}$ is a prime induced subgraph of $G$ and hence is in $\mathcal{H}$. Since $G^{*}\in\mathcal{H}$, from the hypothesis, every blowup of $G^{*}$ is recolorable. Let $H^{*}$ = $G(Q_{1},\dots,Q_{m})$ be the clique skeleton of $G$. Note that $H^{*}$ is a blowup of $G^{*}$ and hence is recolorable. By the vertex minimality of $G$, every proper induced subgraph of $G$ is recolorable. Therefore, since $H^{*}$ is recolorable, by Theorem 6, $G$ is recolorable, a contradiction. ∎
Let $\mathcal{G}$ be a hereditary class of graphs and let $\mathcal{H}$ be the set of prime graphs in $\mathcal{G}$. It follows from Theorem 1 that to determine if every $G\in\mathcal{G}$ is recolorable, it is sufficient to determine if every blowup of every $H\in\mathcal{H}$ is recolorable. We can further ask: to determine if every $G\in\mathcal{G}$ is recolorable, is it sufficient to determine if every $H\in\mathcal{H}$ is recolorable? We answer this in the positive for $2K_{2}$-free graphs.
@@
Lemma 7.
([1]) Let $G$ contain a vertex $v$ of degree at most $\chi(G)$-1. Then, either
@@ 4 $P_{5}$-free graphs
Lemma 8.
([2]) Suppose $G$ has a tight clique cutset. If every proper induced subgraph of $G$ is recolorable, then $G$ is recolorable.
@@
Theorem 9.
([7]) If $G$ is a prime ($P_{5}$, diamond)-free graph containing a $2K_{2}$, then $G$ is a matched co-bipartite graph.
Theorem 10.
([1]) Every ($K_{2}$, diamond)-free graph is recolorable.
@@
Lemma 12.
([14]) Let $G$ = ($V$, $E$) be a prime graph on $n$ vertices. Then $G$ is bipartite and $P_{5}$-free if and only if the following conditions hold:
@@
Theorem 11.
([12]) If $G$ is a prime ($P_{5}$, house, bull)-free graph on 6 or more vertices without a universal vertex, then $G$ is either a bipartite graph or a co-bipartite graph.
@@
Theorem 12.
([13]) Let $G$ be a prime semi-$P_{4}$-sparse graph. Then either $G$ or its complement is a bipartite graph or a thin spider.
@@
Theorem 13.
([14]) Every prime ($P_{5}$, house)-free graph is either a $C_{5}$ or $C_{5}$-free.

### v2

@@ Abstract
The reconfiguration graph of the $k$-colorings of a graph $G$, denoted $R_{k}(G)$, is the graph whose vertices are the $k$-colorings of $G$ and two colorings are adjacent in $R_{k}(G)$ if they differ in color on exactly one vertex. A graph $G$ is said to be recolorable if $R_{\ell}(G)$ is connected for all $\ell\geq\chi(G)$+1. We demonstrate how to use the modular decomposition of a graph class to prove that the graphs in the class are recolorable. In particular, we prove that every ($P_{5}$, diamond)-free graph, every ($P_{5}$, house, bull)-free graph, and every ($P_{5}$, $C_{5}$, co-fork)-free graph is recolorable.
A graph is prime if it cannot be decomposed by modular decomposition except into single vertices. For a prime graph $H$, we study the complexity of deciding if $H$ is $k$-colorable and the complexity of deciding if there exists a path between two given $k$-colorings in $R_{k}(H)$. Suppose $\mathcal{G}$ is a hereditary class of graphs. We prove that if every blowup of every prime graph in $\mathcal{G}$ is recolorable, then every graph in $\mathcal{G}$ is recolorable.
Keywords: reconfiguration graph, $k$-mixing, modular decomposition, $P_{5}$-free graphs.
@@ 1 Introduction
Let $G$ be a finite simple graph with vertex-set $V(G)$ and edge-set $E(G)$. We assume that $G$ is connected unless stated otherwise. We use $n$ to represent the number of vertices in a graph. Two vertices $u$ and $v$ are adjacent in $G$ if $uv\in E(G)$. For a positive integer $k$, a $k$-coloring of $G$ is a mapping from $V(G)$ to a set of colors $\{1,2,\dots,k\}$ such that no two adjacent vertices receive the same color. We say that $G$ is $k$-colorable if it admits a $k$-coloring, and the chromatic number of $G$, denoted $\chi(G)$, is the minimum number of colors required to color $G$. We say $\chi$-coloring of $G$ instead of $\chi(G)$-coloring of $G$, when appropriate. The reconfiguration graph of the $k$-colorings, denoted $R_{k}(G)$, is the graph whose vertices are the $k$-colorings of $G$ and two colorings are adjacent in $R_{k}(G)$ if they differ in color on exactly one vertex. A graph $G$ is said to be $k$-mixing if $R_{k}(G)$ is connected. Given a reconfiguration graph, we can ask the following questions: Is the reconfiguration graph connected? If so, what is the diameter of the reconfiguration graph? In this paper we focus on the connectivity of the reconfiguration graph of the $k$-colorings. A graph is said to be $k$-mixing if $R_{k}(G)$ is connected. We say a graph $G$ is recolorable if $G$ is $\ell$-mixing for all $\ell\geq\chi(G)$+1.
@@
Deciding whether there exists a path between two colorings in $R_{k}(G)$ is PSPACE-complete for $k>3$ [7], and can be decided in polynomial time for $k\leq 3$ [9]. In 2018, Wrochna proved that the problem remains PSPACE-complete for graph classes with bounded bandwidth, and hence for graph classes with bounded treewidth [20]. This led researchers to study the problem for various restricted graph classes. In [6], Bonamy et. al. introduced the class of $k$-color-dense graphs and proved that every $k$-color-dense graph is recolorable. The class of $k$-color-dense graphs includes the class of $k$-colorable chordal graphs. For a graph $H$, a graph $G$ is $H$-$free$ if no induced subgraph of $G$ is isomorphic to $H$. In [3], it was proved that every $H$-free graph $G$ is recolorable if and only if $H$ is an induced subgraph of $P_{4}$ or $P_{3}$+$P_{1}$. For a collection of graphs $\mathcal{H}$, $G$ is $\mathcal{H}$-free if $G$ is $H$-free for every $H\in\mathcal{H}$. For more results about reconfiguration of vertex colorings in forbidden induced subgraph classes see [1].
@@
Bonamy and Bousquet [5] used the tree decomposition of graphs to show that every graph $G$ is ($tw(G)$+2)-mixing, where $tw(G)$ denotes the treewidth of the graph $G$. In this paper we demonstrate how to use the modular decomposition of a graph $G$ to prove that $G$ is $\ell$-mixing for $\ell\geq\chi(G)$+1.
@@
A non-empty set $S\subseteq V(G)$ is a module if every vertex in $V(G)\setminus S$ is either adjacent to every vertex of $S$ or no vertex of $S$. A module $S\subset V(G)$ is said to be non-trivial if $S$ contains at least two vertices. Note that for each $v\in V(G)$, $\{v\}$ is a trivial module in $G$. A graph $G$ is prime if it does not contain any non-trivial module.
@@
A class of graphs $\mathcal{G}$ is said to be hereditary if for every $G$ in $\mathcal{G}$, every induced subgraph of $G$ is also in $\mathcal{G}$. Note that for a collection of graphs $\mathcal{H}$, the class of $\mathcal{H}$-free graphs is hereditary. In Section 3, for a prime graph $H$, we study the complexity of deciding if $H$ is $k$-colorable and the complexity of deciding if there exists a path between two given $k$-colorings in $R_{k}(H)$.
@@
Given a prime graph $H$, for all $k\geq 3$, deciding whether $H$ is $k$-colorable is NP-complete.
Theorem 2.
Given a prime graph $H$, for all $k\geq 4$, deciding whether there exists a path between two given colorings in $R_{k}(H)$ is PSPACE-complete.
In Section 4, we will demonstrate how to use modular decomposition in reconfiguration of vertex coloring.
Theorem 3.
Let $\mathcal{G}$ be a hereditary class of graphs. If every blowup of every prime graph in $\mathcal{G}$ is recolorable, then every graph in $\mathcal{G}$ is recolorable.
We use the above result in Section 5 along with the modular decomposition of several subclasses of $P_{5}$-free graphs to prove the following. (See Figure 1 for graphs mentioned in Theorems 4 and 5.)
@@ 3 Complexity
For a graph $G$, the sibling of $G$ is the graph obtained by adding, for each $x\in V(G)$, a vertex $x^{{}^{\prime}}$ adjacent to only $x$.
Lemma 1.
For any connected graph $G$ the sibling of $G$ is prime.
Proof.
Let $H$ be the sibling of a connected graph $G$ obtained by adding, for each $x\in V(G)$, a vertex $x^{{}^{\prime}}$ adjacent to only $x$. Assume $H$ is not prime. Then $H$ contains a non-trivial module $S^{{}^{\prime}}$.
Suppose $S^{{}^{\prime}}\cap V(G)$ = $\emptyset$. Let $x^{{}^{\prime}}$ and $y^{{}^{\prime}}$ be in $S^{{}^{\prime}}$. Since $x\in N_{H}(x^{{}^{\prime}})\setminus N_{H}(y^{{}^{\prime}})$, $x$ must be in $S^{{}^{\prime}}$, a contradiction. Thus $S^{{}^{\prime}}\cap V(G)\neq\emptyset$.
Claim 1: For any $x\in S^{{}^{\prime}}\cap V(G)$, $x^{{}^{\prime}}\in S^{{}^{\prime}}$.
Let $x\in S^{{}^{\prime}}\cap V(G)$. Suppose $x^{{}^{\prime}}\notin S^{{}^{\prime}}$. Since $S^{{}^{\prime}}$ is a non-trivial module, it must contain some vertex $y$ of $H$ different from $x$. Then $x^{{}^{\prime}}$ is adjacent to $x$ but not to $y$, this contradicts that $S^{{}^{\prime}}$ is a module. This proves Claim 1.
Claim 2: For any $x\in S^{{}^{\prime}}\cap V(G)$, $N_{G}(x)\subseteq S^{{}^{\prime}}$.
Let $x\in S^{{}^{\prime}}\cap V(G)$. By Claim 1, $x^{{}^{\prime}}\in S^{{}^{\prime}}\cap V(H)$. Since every vertex in $N_{G}(x)$ is adjacent to $x$ but not to $x^{{}^{\prime}}$, we have $N_{G}(x)\subseteq S^{{}^{\prime}}$. This proves Claim 2.
Since $H$ and $G$ are connected and $V(H)=V(G)\cup\{x^{{}^{\prime}}\mid x\in V(G)\}$, by Claims 1 and 2, we have $S^{{}^{\prime}}=V(H)$. This contradicts that $S^{{}^{\prime}}$ is a non-trivial module. ∎
Theorem 8 ([16]).
For all $k\geq 3$, deciding whether a connected graph is $k$-colorable is NP-complete.
We are now ready to prove Theorem 1.
Proof of Theorem 1.
Clearly the problem is in NP. To prove that it is NP-complete we will give a reduction from a connected graph to a prime graph. Let $G$ be a connected graph and let $H$ be the sibling of $G$ obtained by adding, for each $x\in V(G)$, a vertex $x^{{}^{\prime}}$ adjacent to only $x$. Note that, for all $k\geq 3$, $H$ is $k$-colorable if and only if $G$ is $k$-colorable. By Lemma 1, $H$ is a prime graph. This proves the theorem. ∎
Theorem 9 ([7]).
Given a connected graph $G$, for all $k\geq 4$, deciding whether there exists a path between two colorings in $R_{k}(G)$ is PSPACE-complete.
We will now prove Theorem 2.
Proof of Theorem 2.
Clearly the problem is in PSPACE. To prove that it is PSPACE-complete we will give a reduction from a connected graph to a prime graph. Let $G$ be a connected graph and let $H$ be the sibling of $G$ obtained by adding, for each $x\in V(G)$, a vertex $x^{{}^{\prime}}$ adjacent to only $x$. Then, by Lemma 1, $H$ is a prime graph. Let $k\geq 4$. Let $\alpha$ and $\beta$ be any two $k$-colorings of $G$. Let $\alpha^{{}^{\prime}}$ and $\beta^{{}^{\prime}}$ be two $k$-colorings of $H$ obtained by extending $\alpha$ and $\beta$, respectively, by coloring each vertex $x^{{}^{\prime}}\in V(H)\setminus V(G)$ a color different from the color of $x$.
Since $G$ is an induced subgraph of $H$, any $k$-coloring of $H$ restricted to $V(G)$ is a $k$-coloring of $G$. Thus if there is a path between $\alpha^{{}^{\prime}}$ and $\beta^{{}^{\prime}}$ in $R_{k}(H)$, then there is a path between $\alpha$ and $\beta$ in $R_{k}(G)$.
Assume that there is a path $P$ between $\alpha$ and $\beta$ in $R_{k}(G)$. Let $\gamma$ and $\psi$ be any two $k$-colorings of $G$ adjacent in the path $P$. Let $\gamma^{{}^{\prime}}$ and $\psi^{{}^{\prime}}$ be any two $k$-colorings of $H$ whose restriction to $V(G)$ is $\gamma$ and $\psi$, respectively. We will prove that there is a path between $\gamma^{{}^{\prime}}$ and $\psi{{}^{\prime}}$ in $R_{k}(H)$. Since $\gamma$ and $\psi$ are adjacent in the path $P$, there exists a unique vertex $y\in V(G)$ such that $\gamma(y)\neq\psi(y)$. Let $y^{{}^{\prime}}\in V(H)\setminus V(G)$ be adjacent to only $y$. Starting from $\gamma^{{}^{\prime}}$, recolor $y^{{}^{\prime}}$ with a color not in $\{\gamma(y),\psi(y)\}$. Recolor the vertex $y$ with color $\psi(y)$. Recolor each vertex $x^{{}^{\prime}}$ in $V(G)\setminus V(H)$, one at a time, with color $\psi^{{}^{\prime}}(x^{{}^{\prime}})$ to obtain the coloring $\psi^{{}^{\prime}}$.
Hence there is a path between $\alpha$ and $\beta$ in $R_{k}(G)$ if and only if there is a path between $\alpha^{{}^{\prime}}$ and $\beta^{{}^{\prime}}$ in $R_{k}(H)$. This completes the proof. ∎
4 Recoloring and Modular Decomposition
@@
For the remainder of this section, we use the following decomposition. Let $G$ be a graph which is neither a join nor a disjoint union of two graphs. Let $S_{1}$,…, $S_{m}$ be a partition of $V(G)$ into maximal modules. Note that $S_{p}\cap S_{q}$ = $\emptyset$ for all $p\neq q$, $p$ and $q\in[m]$. Let $\chi(G)=k$ and let $\chi(G[S_{p}])=k_{p}$, for all $p\in[m]$. For a graph $G$, the clique skeleton of $G$ is the graph obtained from $G$ by recursively substituting a clique $Q_{p}$ of size $k_{p}$ for each $S_{p}$ in $G$, where $p\in[m]$. We use $G(Q_{1},\dots,Q_{m})$ to denote the clique skeleton of $G$. Similar operations have been defined in the literature, for example see [17]. In Figure 2, for a graph $G$, we have illustrated its skeleton and its clique skeleton.
Let $H$ = $G(Q_{1},\dots,Q_{m})$ be the clique skeleton of $G$. Then for any $k$-coloring $\alpha$ of $G$, there exists a $k$-coloring $\beta$ of $H$ such that $\beta(Q_{p})\subseteq\alpha(S_{p})$, for all $p\in[m]$. Note that such a coloring $\beta$ of $H$ can be obtained by coloring each $Q_{p}$ with colors from the set $\alpha(S_{p})$, for all $p\in[m]$. Similarly, given any $k$-coloring $\beta$ of $H$, there exists a $k$-coloring $\alpha$ of $G$ such that for all $p\in[m]$, $\alpha(S_{p})$ = $\beta(Q_{p})$. Thus $\chi(G)$ = $\chi(H)$. Note that the clique skeleton of a graph $G$ is a blowup of the skeleton of $G$.
$a_{1}$$a_{2}$$b_{1}$$b_{2}$$c_{1}$$d_{1}$$e_{1}$$e_{2}$$e_{3}$$e_{4}$$e_{5}$
(a) A graph $G$
$A$$B$$C$$D$$E$
(b) The skeleton of $G$
[…]
(c) The clique skeleton of $G$
Figure 2: A graph $G$, its skeleton, and its clique skeleton.
@@
Let $\mathcal{G}$ be a hereditary class of graphs and let $\mathcal{H}$ be the set of prime graphs in $\mathcal{G}$. Assume that for all $H\in\mathcal{H}$, every blowup of $H$ is recolorable. If possible, let $G\in\mathcal{G}$ be a vertex-minimal graph that is not recolorable. Then, by Lemma 6, $G$ is neither a join nor a disjoint union of two graphs. Let $S_{1}$,…, $S_{m}$ be a partition of $V(G)$ into maximal modules. Let $G^{*}$ be the skeleton of $G$. Then $G^{*}$ is a prime induced subgraph of $G$ and hence is in $\mathcal{H}$. Since $G^{*}\in\mathcal{H}$, from the hypothesis, every blowup of $G^{*}$ is recolorable. Let $H^{*}$ = $G(Q_{1},\dots,Q_{m})$ be the clique skeleton of $G$. Note that $H^{*}$ is a blowup of $G^{*}$ and hence is recolorable. By the vertex minimality of $G$, every proper induced subgraph of $G$ is recolorable. Therefore, since $H^{*}$ is recolorable, by Theorem 10, $G$ is recolorable, a contradiction. ∎
Let $\mathcal{G}$ be a hereditary class of graphs. Then Theorem 3 states that to determine if every graph in $\mathcal{G}$ is recolorable, it is sufficient to determine if every blowup of every prime graph in $\mathcal{G}$ is recolorable. We can further ask: To determine if every graph in $\mathcal{G}$ is recolorable, is it sufficient to determine if every prime graph in $\mathcal{G}$ is recolorable? We answer this in the positive for $2K_{2}$-free graphs and diamond-free graphs.
@@
Lemma 8 ([2]).
Let $G$ contain a vertex $v$ of degree at most $\chi(G)$-1. Then, either
@@ 5 $P_{5}$-free graphs
Lemma 9 ([3]).
Suppose $G$ has a tight clique cutset. If every proper induced subgraph of $G$ is recolorable, then $G$ is recolorable.
@@
Theorem 13 ([8]).
If $G$ is a prime ($P_{5}$, diamond)-free graph containing a $2K_{2}$, then $G$ is a matched co-bipartite graph.
Theorem 14 ([2]).
Every ($K_{2}$, diamond)-free graph is recolorable.
@@
Lemma 13 ([15]).
Let $G$ = ($V$, $E$) be a prime graph on $n$ vertices. Then $G$ is bipartite and $P_{5}$-free if and only if the following conditions hold:
@@
Theorem 15 ([13]).
If $G$ is a prime ($P_{5}$, house, bull)-free graph on 6 or more vertices without a universal vertex, then $G$ is either a bipartite graph or a co-bipartite graph.
@@
Theorem 16 ([14]).
Let $G$ be a prime semi-$P_{4}$-sparse graph. Then either $G$ or its complement is a bipartite graph or a thin spider.
@@
Theorem 17 ([15]).
Every prime ($P_{5}$, house)-free graph is either a $C_{5}$ or $C_{5}$-free.